# Free Energy Calculations using Smooth Basin Classification

### v1

@@ Acknowledgements.
The authors acknowledge funding from the Research Board of Ghent University (V.V.S) and iBOF-21-085 PERSIST (V.V.S. and T.B.). S.V. and P.D. wish to thank the Research Foundation – Flanders (FWO) for doctoral fellowships (grant nos. 11H6821N and 11O2123N respectively). M.B. acknowledges financial support from the Fund for Scientific Research Flanders and the Excellence of Science (EOS) Project BioFact (EOS ID 30902231). We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call. Part of the simulations were performed on the Luxembourg national supercomputer MeluXina. The authors gratefully acknowledge the LuxProvide teams for their expert support.References
- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications (Elsevier Science, 2023).
- Tuckerman [2010] M. Tuckerman, Statistical Mechanics: Theory and Molecular Simulation, Oxford Graduate Texts (OUP Oxford, 2010).
- Grifoni et al. [2019] E. Grifoni, G.-M. Piccini, and M. Parrinello, Microscopic description of acid–base equilibrium, Proc. Natl. Acad. Sci. U.S.A. 116, 4054 (2019).
- Ahalawat and Mondal [2018] N. Ahalawat and J. Mondal, Assessment and optimization of collective variables for protein conformational landscape: GB1 -hairpin as a case study, J. Chem. Phys. 149, 094101 (2018).
- Niu et al. [2018] H. Niu, P. M. Piaggi, M. Invernizzi, and M. Parrinello, Molecular dynamics simulations of liquid silica crystallization, Proc. Natl. Acad. Sci. U.S.A. 115, 5348 (2018).
- Chen [2021] M. Chen, Collective variable-based enhanced sampling and machine learning, Eur. Phys. J. B 94, 211 (2021).
- Bhakat [2022] S. Bhakat, Collective variable discovery in the age of machine learning: reality, hype and everything in between, RSC Adv. 12, 25010 (2022).
- Elishav et al. [2023] O. Elishav, R. Podgaetsky, O. Meikler, and B. Hirshberg, Collective variables for conformational polymorphism in molecular crystals, J. Phys. Chem. Lett. 14, 971 (2023).
- Bonati et al. [2020] L. Bonati, V. Rizzi, and M. Parrinello, Data-driven collective variables for enhanced sampling, J. Phys. Chem. Lett. 11, 2998 (2020).
- Mendels et al. [2018] D. Mendels, G. Piccini, and M. Parrinello, Collective variables from local fluctuations, J. Phys. Chem. Lett. 9, 2776 (2018).
- Pérez-Hernández et al. [2013] G. Pérez-Hernández, F. Paul, T. Giorgino, G. De Fabritiis, and F. Noé, Identification of slow molecular order parameters for markov model construction, J. Chem. Phys. 139, 015102 (2013).
- Wehmeyer and Noé [2018] C. Wehmeyer and F. Noé, Time-lagged autoencoders: Deep learning of slow collective variables for molecular kinetics, J. Chem. Phys. 148, 241703 (2018).
- Mardt et al. [2018] A. Mardt, L. Pasquali, H. Wu, and F. Noé, Vampnets for deep learning of molecular kinetics, Nat. Commun. 9, 5 (2018).
- Thomas et al. [2018] N. Thomas, T. Smidt, S. Kearnes, L. Yang, L. Li, K. Kohlhoff, and P. Riley, Tensor field networks: Rotation- and translation-equivariant neural networks for 3d point clouds (2018), arXiv:1802.08219 [cs.LG] .
- Batzner et al. [2022] S. Batzner, A. Musaelian, L. Sun, M. Geiger, J. P. Mailoa, M. Kornbluth, N. Molinari, T. E. Smidt, and B. Kozinsky, E(3)-equivariant graph neural networks for data-efficient and accurate interatomic potentials, Nat. Commun. 13, 2453 (2022).
- Kovács et al. [2023] D. P. Kovács, I. Batatia, E. S. Arany, and G. Csányi, Evaluation of the MACE force field architecture: From medicinal chemistry to materials science, J. Chem. Phys. 159, 044118 (2023).
- Bishop [2006] C. M. Bishop, Pattern Recognition and Machine Learning (Information Science and Statistics) (Springer-Verlag, Berlin, Heidelberg, 2006).
- Shirts and Chodera [2008] M. R. Shirts and J. D. Chodera, Statistically optimal analysis of samples from multiple equilibrium states, J. Chem. Phys. 129, 124105 (2008).
- Shirts [2017] M. R. Shirts, Reweighting from the mixture distribution as a better way to describe the multistate bennett acceptance ratio (2017), arXiv:1704.00891 [cond-mat.stat-mech] .
- Shirts and Ferguson [2020] M. R. Shirts and A. L. Ferguson, Statistically optimal continuous free energy surfaces from biased simulations and multistate reweighting, J. Chem. Theory Comput. 16, 4107 (2020).
- Shirts et al. [2023] M. Shirts, K. Beauchamp, L. Naden, J. Chodera, jaimergp, S. Martiniani, C. Stern, M. Henry, J. Fass, R. Gowers, R. T. McGibbon, B. Dice, C. Jones, D. L. Dotson, F. York, I. Pulido, M. Thompson, and T. Burgin, choderalab/pymbar: 4.0.2 optional jax build (2023).
- Trizio and Parrinello [2021] E. Trizio and M. Parrinello, From enhanced sampling to reaction profiles, The Journal of Physical Chemistry Letters 12, 8621 (2021).
- Serre et al. [2002] C. Serre, F. Millange, C. Thouvenot, M. Noguès, G. Marsolier, D. Louër, and G. Férey, Very large breathing effect in the first nanoporous chromium(iii)-based solids: MIL-53 or CrIII(OH)·{O2C-C6H4-CO2}·{HO2C-C6H4-CO2H}x·H2Oy, Journal of the American Chemical Society 124, 13519 (2002).
- Demuynck et al. [2018] R. Demuynck, J. Wieme, S. M. J. Rogge, K. D. Dedecker, L. Vanduyfhuys, M. Waroquier, and V. Van Speybroeck, Protocol for identifying accurate collective variables in enhanced molecular dynamics simulations for the description of structural transformations in flexible metal–organic frameworks, Journal of Chemical Theory and Computation 14, 5511 (2018).
- Tan et al. [2024] A. R. Tan, J. C. B. Dietschreit, and R. Gomez-Bombarelli, Enhanced sampling of robust molecular datasets with uncertainty-based collective variables (2024), arXiv:2402.03753 [cs.LG] .
- Bailleul et al. [2020] S. Bailleul, K. Dedecker, P. Cnudde, L. Vanduyfhuys, M. Waroquier, and V. Van Speybroeck, Ab initio enhanced sampling kinetic study on mto ethene methylation reaction, Journal of Catalysis 388, 38 (2020).
- Belkacemi et al. [2022] Z. Belkacemi, P. Gkeka, T. Lelièvre, and G. Stoltz, Chasing collective variables using autoencoders and biased trajectories, J. Chem. Theory Comput. 18, 59 (2022).
- Odstrcil et al. [2022] R. E. Odstrcil, P. Dutta, and J. Liu, Lines: Log-probability estimation via invertible neural networks for enhanced sampling, J. Chem. Theory Comput. 18, 6297 (2022).
- Vlachas et al. [2022] P. R. Vlachas, J. Zavadlav, M. Praprotnik, and P. Koumoutsakos, Accelerated simulations of molecular systems through learning of effective dynamics, J. Chem. Theory Comput. 18, 538 (2022).
- Sun et al. [2022] L. Sun, J. Vandermause, S. Batzner, Y. Xie, D. Clark, W. Chen, and B. Kozinsky, Multitask machine learning of collective variables for enhanced sampling of rare events, J. Chem. Theory Comput. 18, 2341 (2022).
- Bonati et al. [2021] L. Bonati, G. Piccini, and M. Parrinello, Deep learning the slow modes for rare events sampling, Proc. Natl. Acad. Sci. U.S.A. 118, e2113533118 (2021).
- Müllender et al. [2023] L. Müllender, A. Rizzi, M. Parrinello, P. Carloni, and D. Mandelli, Effective data-driven collective variables for free energy calculations from metadynamics of paths (2023), arXiv:2311.05571 [physics.comp-ph] .
- Vandenhaute et al. [2023] S. Vandenhaute, M. Cools-Ceuppens, S. DeKeyser, T. Verstraelen, and V. Van Speybroeck, Machine learning potentials for metal-organic frameworks using an incremental learning approach, npj Comput. Mater. 9, 19 (2023).
- Liao and Smidt [2023] Y.-L. Liao and T. Smidt, Equiformer: Equivariant graph attention transformer for 3d atomistic graphs (2023), arXiv:2206.11990 [cs.LG] .
- Pelaez et al. [2024] R. P. Pelaez, G. Simeon, R. Galvelis, A. Mirarchi, P. Eastman, S. Doerr, P. Thölke, T. E. Markland, and G. D. Fabritiis, Torchmd-net 2.0: Fast neural network potentials for molecular simulations (2024), arXiv:2402.17660 [cs.LG] .
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills, and G. Csányi, A foundation model for atomistic materials chemistry (2023), arXiv:2401.00096 [physics.chem-ph] .
- Deng et al. [2023] B. Deng, P. Zhong, K. Jun, J. Riebesell, K. Han, C. J. Bartel, and G. Ceder, Chgnet as a pretrained universal neural network potential for charge-informed atomistic modelling, Nature Machine Intelligence 5, 1031 (2023).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).

### v3

@@ ACKNOWLEDGEMENTS
The authors acknowledge funding from the Research Board of Ghent University (V.V.S) and iBOF-21-085 PERSIST (V.V.S. and T.B.). S.V. and P.D. wish to thank the Research Foundation – Flanders (FWO) for doctoral fellowships (grant nos. 11H6821N and 11O2123N respectively). M.B. acknowledges financial support from the Fund for Scientific Research Flanders and the Excellence of Science (EOS) Project BioFact (EOS ID 30902231). We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call. Part of the simulations were performed on the Luxembourg national supercomputer MeluXina. The authors gratefully acknowledge the LuxProvide teams for their expert support.
Appendix A Appendix on the Model Architecture
In analogy with the prediction of the total potential energy as a sum of predicted atomic energies , we predict the total log probability as a sum of per-atom log probabilities (see Equation 2 in the main text). In a system with basins, the log probabilities are a vector of length , i.e. there is one per-atom logit for each basin. We use to denote the GNN features from layer for atom . First, we forward the node features into a MLP per layer:
Next, we concatenate for all GNN layers , and use a final linear layer with weights to compute the per-atom log probabilities :
Finally, the total log probabilities are obtained by summing the individual
The phase readout layers are analogous to the existing energy readouts found in conventional GNNs [20, 21, 22] (Figure 1, c), and essentially consist of a multilayer perceptron (MLP) containing at most two nonlinearities and about 1000 weights in total (compared to 100,000 weights in the GNN).
Appendix B Appendix on Model Training
The predicted log probabilities are normalized and trained to atomic geometries which are labeled with the correct basin. The label is obtained by minimizing the energy of the structure, as explained in Figure 1 in the main manuscript. In a classification setting, one often uses the cross-entropy loss function [24]. We can define the phase label of an atomic geometry as a one-hot vector which has the same length as (i.e. the number of basins), and whose components indicate the correct basin of attraction:
Based on Equation 8, we can write the cross-entropy loss as a simple scalar product:
The classifier is trained while keeping the GNN and energy readouts fixed, similar to conventional transfer learning approaches. Its training data consists of atomic geometries which are labeled with their corresponding basin of attraction (e.g. A or B).
Note that the cross-entropy loss is a function defined on the total log probabilities , not on the specific per-atom log probabilities . This implies that the cross-entropy loss will incentivize the model to predict correct total log probabilities, but otherwise introduces no restrictions regarding the specific per-atom log probabilities . This makes the model prone to overfitting, especially in the low-data regime. To improve this, we considered a number of regularization strategies including label smoothing, an L2 penalty on the weights, and an L1 penalty on logit gradients with respect to the atomic coordinates. Their efficacy was highly dependent on both the system under study and their specific weights in the loss function, and we chose not to apply them. Instead, we designed a new regularization term which prevents the per-atom log probabilities from becoming strongly positive for any given atomic environment .
First, note that the global probability for each phase can be computed based on the predicted log probabilities using a softmax function:
for any class A. We can apply a similar reasoning to the per-atom log probabilities in order to obtain per-atom phase probabilities . Next, consider
The quantity denotes an expected per-atom logit value. In practice, given the exponential dependence between the probabilities and the log probabilities, there will only be one nonzero contribution in the sum in Equation 11, and the expectation value will essentially reduce to the largest (per-atom) logit value. Note that the predicted per-atom phase (based on the per-atom log probabilities) can differ from the predicted global phase (based on the total log probabilities). Finally, in the loss function, we add an L2 regularization on , which essentially prevents the model from being overconfident in any of the per-atom log probabilities:
For all experiments in this work, we use .
Furthermore, note that the actual training of the phase readouts can be made very efficient because the GNN node features can be cached and reused at every step of the optimization since the weights of the pretrained GNN are fixed. As a result, training is very efficient, and all SBC models in this work can be trained within one hour on a consumer-level GPU.
Appendix C Appendix on the Active Learning
Figure 1 (e) in the main manuscript outlines the active learning workflow in more detail. We start by training the phase readouts on equilibrium samples from each of the free energy minima, after which we begin with the active learning loop and perform enhanced sampling. We introduce harmonic restraints centered on different values of in order to force the system to transition from one basin to another. In each iteration, we position 32 umbrellas along the entire CV range, at three different temperatures as to ensure that both low- and (relevant) high-energy regions are properly sampled. In each umbrella, the system is randomly initialized in either of the two minima. The samples from the resulting trajectory are labeled according to their basin of attraction, as determined by a simple energy minimization for each sample. After this, the model is retrained on all data, and the procedure is repeated. See the Supplemental Material (Section S1) for a complete overview of all hyperparameters.
Appendix D Appendix on the Graph Neural Network
All calculations in this work employed MACE as base potential [22], although we should emphasize that SBC is independent of the specific GNN architecture [40]. The interatomic potentials for MIL-53(Al) and alanine dipeptide were trained from scratch using psiflow, at the PBE-D3(BJ) level of theory and with a triple zeta basis set [37]. DFT evaluations were performed with CP2K 2023.1 for MIL-53(Al) in which the basis set was augmented with plane waves up until a cutoff of 1000 Ry, and ORCA 5.0.4 for alanine dipeptide [41, 42]. For the proton transfer reaction in the zeolite catalyst, we employed a foundational MACE model [38]. For the MACE models, we employ node features with and 16 channels each, with a cutoff radius of 6.5 Å. All models will be made available upon final publication.
Appendix E Appendix on the Free Energy Calculations
in which we emphasize that the bias energy contains an implicit dependence on both the atomic coordinates and the unit cell parameters through the learned atomic environment features [23]. We use the multistate Bennett acceptance ratio (MBAR) equations to combine samples from multiple umbrellas in a statistically optimal manner and obtain free energy profiles as shown in Figure 4 [43, 44, 45, 46]. Note that while relative stabilities are intrinsic to the system, the actual free energy profiles depend on the particular choice of CV, so the learned and reference free energy profiles do not need to have the same shape (although there exist ways to transform one profile into another [47]).